OpenStudy (anonymous):

Tell the property used in each of the steps. One point each for each correct property. Steps Property 3 + 4(x + 1) Given = 3 + (4x + 4) A = (4x + 4) + 3 B = 4x + (4 + 3) C = 4x + 7 D

OpenStudy (anonymous):

I know A is distributive property, but i dont know the rest, please help!!

OpenStudy (anonymous):

medal for anyone who can help me

OpenStudy (ajprincess):

a+b=b+a=>commutative law is this similar to either B or C?

OpenStudy (anonymous):

its similar to B, and A

OpenStudy (ajprincess):

so what is the property described by B?

OpenStudy (anonymous):

commutative property of addition

OpenStudy (ajprincess):

yup right:)

OpenStudy (ajprincess):

a+(b+c)=(a+b)+c=>associative law

OpenStudy (anonymous):

So im guessing c would be associative because the only thing that moves is the parentheses

OpenStudy (ajprincess):

yup it is associative:)

OpenStudy (anonymous):

What would D be though, because that's just addition

OpenStudy (ajprincess):

am really sorry. no idea abt D

OpenStudy (anonymous):

Lol, no problem your a life saver though, Thanks!!